Device Status
OLT : OK
UPS: OK
ONU: Ok
POE Switch : OK
Findings: Too much dust inside all device.
Suggestion: New separate panel Installation for OLT device.
Device Status
OLT : OK
UPS: OK
ONU: Ok
POE Switch : OK
Findings: Too much dust inside all device.
Suggestion: New separate panel Installation for OLT device.
Problem: When the pump is stopped and normally shuts off, a fault automatically occurs in the MOV (4P01-E). In this case the valve…
DTW NO. LIT PIT Comments DTW NO. LIT PIT Comments 101 NOT OK OK 201 OK OK 102 OK OK 202 NOT OK…
Observation: DTW 107 Radio Communication checking complete. This DTWs communication is very bad.1. In DTW 107, We checked up to 10 minutes. In…
Due to solving those problems : Modbus ground connection check. VFD frequency regulator loose connection check. SPD check . RJ-45 connector check. Cable…
Due To Strom the optical fiber cable fall down and damage near DTW-209 to 206 Now Everything is Normal
1.Cable connector change 2. IP Setup
One Comment